### v1

@@ Abstract
Context. The James Webb Space Telescope (JWST) has revealed a population of active galactic nuclei (AGNs) that challenge existing black hole (BH) formation models. These newly observed BHs are overmassive compared to the host galaxies and have an unexpectedly high abundance. Their exact origin remains elusive.
@@ 1 Introduction
Recent JWST surveys have unveiled a large population of faint AGN candidates at $z\geq$ 4 (Kokorev et al. 2023; Furtak et al. 2024; Lambrides et al. 2024; Larson et al. 2023; Goulding et al. 2023; Maiolino et al. 2024; Bogdán et al. 2024; Kovács et al. 2024; Matthee et al. 2024; Kokorev et al. 2024; Greene et al. 2024). Among the most intriguing discoveries are very compact, reddish objects that exhibit broad line emission features in their spectra, known as the little red dots (LRDs; Matthee et al. 2024; Greene et al. 2024; Kocevski et al. 2024; Maiolino et al. 2025). These newly discovered BHs have masses of a few times $10^{5}-10^{7}$ M, less massive than quasars at the same redshifts but, interestingly, similar in mass at the low end to those expected for direct collapse black holes at birth (DCBHs; Hosokawa et al. 2013; Woods et al. 2017; Latif et al. 2022b; Patrick et al. 2023). They are overmassive relative to their host galaxies when compared to the local $\rm M_{BH}-M_{star}$ relation and their bolometric luminosities range from $\rm 10^{44}-10^{45}\leavevmode\nobreak\ erg/s$. They reside in galaxies with sub-solar metallicities. These discoveries provide key insights into the AGN population, the co-evolution of AGNs with their host galaxies, BH seeding, and growth mechanisms. The cosmic abundance of these recently observed AGN is $\rm\sim 10^{-5}\ Mpc^{-3}\ mag^{-1}$, about two orders of magnitude higher than the faintest UV selected quasars observed at such redshifts (Matthee et al. 2024; Kokorev et al. 2024; Greene et al. 2024).
@@
On the theoretical side, several possibilities exist to explain the observed SMBHs. The most straightforward path for BH seed formation in the early Universe is the gravitational collapse of massive Pop III stars. Despite the uncertainties in the Pop III IMF, the general consensus is to consider BH seeds of 100 M. If the $z\gtrsim 7$ quasars are descendants of these light seeds, they must have grown uninterruptedly at the Eddington rate. Even for some newly discovered JWST AGNs, super-Eddington accretion would be necessary for such low initial BH masses (Bogdán et al. 2024; Kovács et al. 2024; Lupi et al. 2024). However, numerical simulations have shown that the birthplaces of these BHs are rapidly depleted of gas due to supernova (SN) feedback, that these BH seeds tend to wander within their host galaxies due to inefficient dynamical friction, and that they rarely encounter cold, dense gas from which to accrete mass. Moreover, they must compete with star formation for the available cold gas (Johnson & Bromm 2007; Alvarez et al. 2009; Pacucci et al. 2017; Smith et al. 2018; Latif & Khochfar 2020). Natural explanations for the observation of supermassive black holes at $z>9$ have been put forward e.g. by Schneider et al. (2023).
@@
Some of the hurdles faced by this formation channel are easily overcome by the DCBH scenario, which produces BH seeds of $10^{4}-10^{5}$ M, born in gas-rich environments (see e.g. Regan & Haehnelt 2009a, b; Latif et al. 2013, 2016; Chon et al. 2018; Wise et al. 2019; Regan et al. 2020; Chon & Omukai 2020; Latif et al. 2022b; Schleicher et al. 2023; Saavedra-Bastidas et al. 2024). However, many uncertainties remain regarding the abundance expected from the DCBH scenario, with estimates differing by several orders of magnitude (Habouzit et al. 2016). It now seems that even optimistic assumptions about DCBH formation are incapable of accounting for the high abundance of high-$z$ AGNs, assuming that all LRDs are indeed powered by SMBHs. Recent models seem to favor lighter but more abundant seeds (Latif et al. 2021; Sassano et al. 2021; Trinca et al. 2024; Regan & Volonteri 2024; O’Brennan et al. 2025).
@@
In light of this new evidence, the exploration of alternative BH formation pathways is worth pursuing. One such mechanism consists of the growth of a central object due to stellar collisions in very dense star clusters (Omukai et al. 2008; Katz et al. 2015; Sakurai et al. 2017, 2019; Reinoso et al. 2018, 2020; Vergara et al. 2021; Rantala et al. 2024). This formation channel has the advantage of producing more massive seeds than the Pop III remnants, and they are embedded in massive systems which aid during the sinking to the galactic center (Mukherjee et al. 2024). Moreover, these seeds should be more abundant than DCBH seeds, which would alleviate tensions with recent JWST data. Semi-analytic models show that they could contribute quite substantially to explain the SMBH population in the present-day Universe (Liempi et al. 2025). Interestingly, the dense stellar systems in which these seeds should form, with effective radii $\lesssim 1$ pc and masses $M\gtrsim 10^{6}$ M, are now being revealed by JWST observations (Adamo et al. 2024). An alternative pathway, in case the stellar evolution is faster than the dynamical evolution of the cluster, consists of the formation of a cluster of stellar mass BHs which may migrate to the center and lead to the formation of a massive BH via gravitational wave emission if the potential of the BH clusters is sufficiently steep, for example due to an external gas inflow (Davies et al. 2011; Lupi et al. 2014; Kroupa et al. 2020; Gaete et al. 2024).
@@ 2.1 Cosmological simulation setup
Numerical simulations are conducted with the cosmological hydrodynamics open source code ENZO (Bryan et al. 2014) which uses an adaptive mesh refinement approach to the cover a large range of spatial scales: from tens of Mpc down to sub parsec scales. ENZO employs a 3rd order piece wise parabolic solver for hydrodynamics, a $N$-body particle-mesh to compute DM dynamics and multigrid Poisson solver for self-gravity calculations. Our simulations start at $z$=200 with cosmological initial conditions generated from Gaussian random density perturbations with MUSIC (Hahn & Abel 2011) and use the following cosmological parameters from the Planck data: $\Omega_{\mathrm{M}}=$ 0.3089, $\Omega_{\Lambda}=$ 0.691, $\Omega_{\mathrm{b}}=$ 0.0486, $h=$ 0.677, $\sigma_{8}=$ 0.816, and $n=$ 0.967 (Planck Collaboration et al. 2016). They are performed in a volume of $\rm(50\leavevmode\nobreak\ Mpc/h)^{3}$ and were initially ran down to $z$=7 with coarse DM resolution of $\rm 512^{3}$ which yields a DM mass resolution of $\rm 10^{8}$ M, sufficient enough to resolve the target halo of $\rm 1.6\times 10^{12}$ M at $z$=7 by 10,000 DM particles. Using the rockstar halo finder (Behroozi et al. 2013), we identify the DM particles belonging to the halo in its Lagrange volume and trace them back to the initial redshift. We then add one refinement level in a region twice the Lagrange radius of the halo and run a zoom-in simulation. We keep repeating these steps, every time adding one additional refinement level in the region twice the Lagrange volume of the target halo until we achieve a DM resolution of a few thousand solar masses ($\sim$ 3600 M) in the zoomed in region111see https://github.com/jwise77/enzo-mrp-music for details. This approach allows us to fully resolve minihalos of $\rm 10^{5}$ M. We further employ up to 18 dynamical levels of refinement in the region of interest during the course of simulation, which enable us to resolve the gravitational collapse down to the scales of about 0.01 pc. Our refinement criteria are based on the baryonic overdensity, DM mass resolution and the Jeans refinement, for details see (Latif et al. 2022a; Latif & Khochfar 2020; Latif et al. 2022b). We resolve the Jeans length by at least 64 cells which is found to be sufficient to resolve turbulent eddies in cosmological simulations (Federrath et al. 2011; Latif et al. 2013). Similar to Latif et al. (2022b), cold flows drive supersonic turbulence that prevents star formation in the halo until it reaches a mass of a few times $\rm 10^{7}\leavevmode\nobreak\ M_{\odot}$ at z=30 and catastrophic gravitational collapse gets triggered.
@@
At this stage, we turn on sink particles which represent Pop III stars in our simulations and are typically formed at densities of $10^{-18}\rm g/cm^{3}$. Our recipe for sink formation is based on SmartStars (Regan & Downes 2018; Latif et al. 2022a) and sinks are created in a grid cell when it meets the following criteria: I) at the maximum refinement level, II) at a local minimum of the gravitational potential, III) has a convergent flow, IV) density is greater than the Jeans density, V) the gas cooling time is shorter than the collapse time. A sink particle can accrete and gets merged with more massive sink forming in the accretion radius corresponding to 4 cells. The stellar feedback is modeled assuming a black-body spectrum with a temperature of $\rm T_{eff}=$ 5000 K for $\rm M_{*}\leq$ 10 M (Hosokawa et al. 2013) and $\rm T_{eff}=10^{4.759}$ when $\rm M_{*}>$ 10 M (Schaerer 2002). The radiative transfer scheme, MORAY, is used to trace radiation from stars. The flux is divided into five energy bins: two for dissociating photons (2.0 eV, 12.8 eV) and three for ionizing photons (14.0 eV, 25.0 eV, 200.0 eV), with bin definitions determined using the SEDOP code developed by Mirocha et al. (2012). For further details, the reader is referred to Latif et al. (2022a). MORAY is coupled to hydrodynamics and non-equilibrium chemistry solver based on Abel et al. (1997). The chemical model solves the rate equations for nine primordial species: $\rm H,\leavevmode\nobreak\ H^{+},\leavevmode\nobreak\ H^{-},\leavevmode\nobreak\ He,\leavevmode\nobreak\ He^{+},\leavevmode\nobreak\ He^{++},\leavevmode\nobreak\ H_{2},\leavevmode\nobreak\ H_{2}^{+},\leavevmode\nobreak\ e^{-}$ and and includes all relevant processes, such as $\rm H_{2}$ formation via three-body reactions, $\rm H_{2}$ cooling as well as cooling from collisional excitation, collisional ionization, radiative recombination, Bremsstrahlung radiation, CIE cooling, $\rm H_{2}$ optically thick cooling approximation at densities $\rm\geq 10^{-14}\leavevmode\nobreak\ g/cm^{3}$ and heating from three-body reactions. We take the simulation output from ENZO at $z=39.4$ when a total of 68 sink particles have formed, containing a total stellar mass of 1497 M and feed it to AMUSE. The simulation was evolved for 330 Kyrs and a snapshot of the state of the system at this time is shown in Fig. 1.
@@ 2.4 The background potential
with $r=2$ pc and $\dot{M}_{\rm gas}=2$ M yr-1 taken from the ENZO run. This results in a value of $\dot{\rho}_{0}=0.142$ M yr-1, which is used as the inflow rate for five models. We also explore a case with a larger inflow rate of $\dot{M}_{\rm gas}=0.9$ M yr-1 at $r=1$ pc, resulting in $\dot{\rho}_{0}=0.238$ M yr-1.
@@ 2.5 The IMF
$\alpha_{1}$ $=-0.66,$ $0.4\leq$ $\ \ m/{\rm M_{\odot}}<3.0,$
$\alpha_{2}$ = $=\ \ \ 0.15,$ $3.0\leq$ $\ \ m/{\rm M_{\odot}}<10.0,$
$\alpha_{3}$ $=-0.83,$ $10.0\leq$ $\ \ m/{\rm M_{\odot}}<20.0,$
$\alpha_{4}$ $=-1.75,$ $20.0\leq$ $\ \ m/{\rm M_{\odot}}<150.0.$
@@ 2.6 Star formation
In our models, we maintain a constant value of the star formation efficiency, which we define as
@@
where $M_{\rm gas,2}$ is the gas mass enclosed within the central 2 pc. The amount of gas that is converted into new stars $M_{\rm new\ stars}$ is defined as:
$M_{\rm new\ stars}=\left\{\begin{aligned} &\ \epsilon_{\rm SF}M_{\rm gas,2}-M_{\rm stars}(1-\epsilon_{\rm SF}),&\epsilon_{\rm SF\ now}<\epsilon_{\rm SF},\\ &\ \epsilon_{\rm SF}\dot{M}_{\rm gas}dt_{\rm SF},&\epsilon_{\rm SF\ now}\geq\epsilon_{\rm SF}.\end{aligned}\right.$ (8)
Once the total mass in new stars has been decided, we draw stars with a uniform probability until the required mass has been reached. A safe check is imposed such that if $M_{\rm new\ stars}>M_{\rm gas,2}$, then star formation is postponed to the next timestep.

### v2

@@ Abstract
Context. The James Webb Space Telescope (JWST) has revealed a population of active galactic nuclei (AGNs) that challenge existing black hole (BH) formation models. These newly observed BHs seem overmassive compared to the host galaxies and have an unexpectedly high abundance. Their exact origin remains elusive.
@@ 1 Introduction
Recent JWST surveys have unveiled a large population of faint AGN candidates at $z\geq$ 4 (Kokorev et al. 2023; Furtak et al. 2024; Lambrides et al. 2024a; Larson et al. 2023; Goulding et al. 2023; Maiolino et al. 2024; Bogdán et al. 2024; Kovács et al. 2024; Matthee et al. 2024; Kokorev et al. 2024; Greene et al. 2024; Napolitano et al. 2024). Among the most intriguing discoveries are very compact, reddish objects that exhibit broad line emission features in their spectra, known as the little red dots (LRDs; Matthee et al. 2024; Greene et al. 2024; Kocevski et al. 2024; Maiolino et al. 2025). These newly discovered BHs have masses of a few times $10^{5}-10^{7}$ M, less massive than quasars at the same redshifts but, interestingly, similar in mass at the low end to those expected for direct collapse black holes at birth (DCBHs; Hosokawa et al. 2013; Woods et al. 2017; Latif et al. 2022b; Patrick et al. 2023). They seem overmassive relative to their host galaxies when compared to the local $\rm M_{BH}-M_{star}$ relation (also see Lupi et al. 2024; King 2024; Lambrides et al. 2024b; Rusakov et al. 2025; Naidu et al. 2025)) and their bolometric luminosities range from $\rm 10^{44}-10^{45}\leavevmode\nobreak\ erg/s$. They reside in galaxies with sub-solar metallicities. These discoveries provide key insights into the AGN population, the co-evolution of AGNs with their host galaxies, BH seeding, and growth mechanisms. The cosmic abundance of these recently observed AGN is $\rm\sim 10^{-5}\ Mpc^{-3}\ mag^{-1}$, about two orders of magnitude higher than the faintest UV selected quasars observed at such redshifts (Matthee et al. 2024; Kokorev et al. 2024; Greene et al. 2024).
@@
On the theoretical side, several possibilities exist to explain the observed SMBHs. The most straightforward path for BH seed formation in the early Universe is the gravitational collapse of massive Pop III stars. Despite the uncertainties in the Pop III IMF, the general consensus is to consider BH seeds of 100 M. If the $z\gtrsim 7$ quasars are descendants of these light seeds, they must have grown uninterruptedly at the Eddington rate. Even for some newly discovered JWST AGNs, super-Eddington accretion would be necessary for such low initial BH masses (Bogdán et al. 2024; Kovács et al. 2024; Lupi et al. 2024; Napolitano et al. 2024). However, numerical simulations have shown that the birthplaces of these BHs are rapidly depleted of gas due to supernova (SN) feedback, that these BH seeds tend to wander within their host galaxies due to inefficient dynamical friction, and that they rarely encounter cold, dense gas from which to accrete mass. Moreover, they must compete with star formation for the available cold gas (Johnson & Bromm 2007; Alvarez et al. 2009; Pacucci et al. 2017; Smith et al. 2018; Latif & Khochfar 2020). Natural explanations for the observation of supermassive black holes at $z>9$ have been put forward e.g. by Schneider et al. (2023).
@@
Some of the hurdles faced by this formation channel are easily overcome by the DCBH scenario, which produces BH seeds of $10^{4}-10^{5}$ M, born in gas-rich environments (see e.g. Regan & Haehnelt 2009a, b; Latif et al. 2013, 2016; Chon et al. 2018; Wise et al. 2019; Regan et al. 2020; Chon & Omukai 2020; Latif et al. 2022b; Schleicher et al. 2023; Saavedra-Bastidas et al. 2024; Chon & Omukai 2024). Depending on the strength of the radiative background, the masses can also lie within 103 to 104 M (Latif et al. 2014; Latif & Volonteri 2015; Regan et al. 2020; Latif et al. 2021). Many uncertainties remain regarding the abundance expected from the DCBH scenario, with estimates differing by several orders of magnitude (Habouzit et al. 2016). It now seems that even optimistic assumptions about DCBH formation are incapable of accounting for the high abundance of high-$z$ AGNs, assuming that all LRDs are indeed powered by SMBHs. Recent models seem to favor lighter but more abundant seeds (Latif et al. 2021; Sassano et al. 2021; Trinca et al. 2024; Regan & Volonteri 2024; O’Brennan et al. 2025).
@@
In light of this new evidence, the exploration of alternative BH formation pathways is worth pursuing. One such mechanism consists of the growth of a central object due to stellar collisions in very dense star clusters (Omukai et al. 2008; Katz et al. 2015; Sakurai et al. 2017, 2019; Reinoso et al. 2018, 2020; Vergara et al. 2021; Rantala et al. 2024). This formation channel has the advantage of producing more massive seeds than the Pop III remnants, and they are embedded in massive systems which aid during the sinking to the galactic center (Mukherjee et al. 2025). Moreover, these seeds should be more abundant than DCBH seeds, which would alleviate tensions with recent JWST data. Semi-analytic models show that they could contribute quite substantially to explain the SMBH population in the present-day Universe (Liempi et al. 2025). Interestingly, the dense stellar systems in which these seeds should form, with effective radii $\lesssim 1$ pc and masses $M\gtrsim 10^{6}$ M, are now being revealed by JWST observations (Adamo et al. 2024). An alternative pathway, in case the stellar evolution is faster than the dynamical evolution of the cluster, consists of the formation of a cluster of stellar mass BHs which may migrate to the center and lead to the formation of a massive BH via gravitational wave emission if the potential of the BH clusters is sufficiently steep, for example due to an external gas inflow (Davies et al. 2011; Lupi et al. 2014; Kroupa et al. 2020; Gaete et al. 2024).
@@ 2.1 Cosmological simulation setup
Numerical simulations are conducted with the cosmological hydrodynamics open source code ENZO (Bryan et al. 2014) which uses an adaptive mesh refinement approach to the cover a large range of spatial scales: from tens of Mpc down to sub parsec scales. ENZO employs a 3rd order piece wise parabolic solver for hydrodynamics, a $N$-body particle-mesh to compute DM dynamics and multigrid Poisson solver for self-gravity calculations. Our simulations start at $z$=200 with cosmological initial conditions generated from Gaussian random density perturbations with MUSIC (Hahn & Abel 2011) and use the following cosmological parameters from the Planck data: $\Omega_{\mathrm{M}}=$ 0.3089, $\Omega_{\Lambda}=$ 0.691, $\Omega_{\mathrm{b}}=$ 0.0486, $h=$ 0.677, and $n=$ 0.967 (Planck Collaboration et al. 2016). They are performed in a volume of $\rm(50\leavevmode\nobreak\ Mpc/h)^{3}$ and were initially ran down to $z$=7 with coarse DM resolution of $\rm 512^{3}$ which yields a DM mass resolution of $\rm 10^{8}$ M, sufficient enough to resolve the target halo of $\rm 1.6\times 10^{12}$ M at $z$=7 by 10,000 DM particles. To capture such a rare halo, we boosted the value of $\sigma_{8}$ to 1.2 as this approach has been used in previous works, see Hirano et al. (2017). Using the rockstar halo finder (Behroozi et al. 2013), we identify the DM particles belonging to the halo in its Lagrange volume and trace them back to the initial redshift. We then add one refinement level in a region twice the Lagrange radius of the halo and run a zoom-in simulation. We keep repeating these steps, every time adding one additional refinement level in the region twice the Lagrange volume of the target halo until we achieve a DM resolution of a few thousand solar masses ($\sim$ 3600 M) in the zoomed in region111see https://github.com/jwise77/enzo-mrp-music for details. This approach allows us to resolve minihalos of $\rm 10^{5}$ M. We further employ up to 18 dynamical levels of refinement in the region of interest during the course of simulation, which enable us to resolve the gravitational collapse down to the scales of about 0.01 pc. Our refinement criteria are based on the baryonic overdensity, DM mass resolution and the Jeans refinement, for details see (Latif et al. 2022a; Latif & Khochfar 2020; Latif et al. 2022b). We resolve the Jeans length by at least 64 cells which is found to be sufficient to resolve turbulent eddies in cosmological simulations (Federrath et al. 2011; Latif et al. 2013). Similar to Latif et al. (2022b), cold flows drive supersonic turbulence that prevents star formation in the halo until it reaches a mass of a few times $\rm 10^{7}\leavevmode\nobreak\ M_{\odot}$ at z=30 (see Fig. 12 which shows the comparison of turbulent and radial infall velocites at different times) and catastrophic gravitational collapse gets triggered. Dense cold streams of gas penetrate the halo and drive turbulent motions of a few 10 km/s, as shown in Fig. 11.
@@
At this stage, we turn on sink particles which represent Pop III stars in our simulations and are typically formed at densities of $10^{-18}\rm g/cm^{3}$. Our recipe for sink formation is based on SmartStars (Regan & Downes 2018; Latif et al. 2022a) and sinks are created in a grid cell when it meets the following criteria: I) at the maximum refinement level, II) at a local minimum of the gravitational potential, III) has a convergent flow, IV) density is greater than the Jeans density, V) the gas cooling time is shorter than the collapse time. A sink particle can accrete and gets merged with more massive sink forming in the accretion radius corresponding to 4 cells. The stellar feedback is modeled assuming a black-body spectrum with a temperature of $\rm T_{eff}=$ 5000 K for $\rm M_{*}\leq$ 10 M (Hosokawa et al. 2013) and $\rm T_{eff}=10^{4.759}$ when $\rm M_{*}>$ 10 M (Schaerer 2002). The radiative transfer scheme, MORAY, is used to trace radiation from stars. The flux is divided into five energy bins: two for dissociating photons (2.0 eV, 12.8 eV) and three for ionizing photons (14.0 eV, 25.0 eV, 200.0 eV), with bin definitions determined using the SEDOP code developed by Mirocha et al. (2012). For further details, the reader is referred to Latif et al. (2022a). MORAY is coupled to hydrodynamics and non-equilibrium chemistry solver based on Abel et al. (1997). The chemical model solves the rate equations for nine primordial species: $\rm H,\leavevmode\nobreak\ H^{+},\leavevmode\nobreak\ H^{-},\leavevmode\nobreak\ He,\leavevmode\nobreak\ He^{+},\leavevmode\nobreak\ He^{++},\leavevmode\nobreak\ H_{2},\leavevmode\nobreak\ H_{2}^{+},\leavevmode\nobreak\ e^{-}$ and and includes all relevant processes, such as $\rm H_{2}$ formation via three-body reactions, $\rm H_{2}$ cooling as well as cooling from collisional excitation, collisional ionization, radiative recombination, Bremsstrahlung radiation, CIE cooling, $\rm H_{2}$ optically thick cooling approximation at densities $\rm\geq 10^{-14}\leavevmode\nobreak\ g/cm^{3}$ and heating from three-body reactions. We ignore deuterium and related species because HD cooling mainly occurs in relic H II regions or shock-heated gas during major mergers which boost D+ abundance (Yoshida et al. 2007; Maio et al. 2007; McGreer & Bryan 2008; Greif et al. 2008; Bovino et al. 2014). Therefore, we expect it not to be too relevant at least in massive halos like the one simulated here. We take the simulation output from ENZO at $z=39.4$ when a total of 68 sink particles have formed, containing a total stellar mass of 1497 M and feed it to AMUSE. The simulation was evolved for 330 Kyrs and a snapshot of the state of the system at this time is shown in Fig. 1.
@@ 2.4 The background potential
with $r=2$ pc and $\dot{M}_{\rm gas}=2$ M yr-1 taken from the ENZO run. This results in a value of $\dot{\rho}_{0}=0.142$ M pc-3 yr-1, which is used as the inflow rate for five models. We also explore a case with a larger inflow rate of $\dot{M}_{\rm gas}=0.9$ M yr-1 at $r=1$ pc, resulting in $\dot{\rho}_{0}=0.238$ M pc-3 yr-1.
@@ 2.5 The IMF
$\alpha_{1}$ = $=\ \ \ 0.66,$ $0.4\leq$ $\ \ m/{\rm M_{\odot}}<3.0,$
$\alpha_{2}$ $=-0.15,$ $3.0\leq$ $\ \ m/{\rm M_{\odot}}<10.0,$
$\alpha_{3}$ = $=\ \ \ 0.83,$ $10.0\leq$ $\ \ m/{\rm M_{\odot}}<20.0,$
$\alpha_{4}$ = $=\ \ \ 1.75,$ $20.0\leq$ $\ \ m/{\rm M_{\odot}}<150.0.$
@@ 2.6 Star formation
During the simulations we define the star formation efficiency as
@@
where $M_{\rm gas,2}$ is the gas mass enclosed within the central 2 pc, and $M_{\rm stars}$ is the current stellar mass. The amount of gas that is converted into new stars $M_{\rm new\ stars}$ is defined as:
$M_{\rm new\ stars}=\left\{\begin{aligned} &\ \epsilon_{\rm SF}M_{\rm gas,2}-M_{\rm stars}(1-\epsilon_{\rm SF}),&\epsilon_{\rm SF\ now}<\epsilon_{\rm SF},\\ &\ \epsilon_{\rm SF}\dot{M}_{\rm gas}dt_{\rm SF},&\epsilon_{\rm SF\ now}\geq\epsilon_{\rm SF},\end{aligned}\right.$ (8)
where $\epsilon_{\rm SF}$ is the SF efficiency assumed for our model, as presented in Table 1. The use of $\epsilon_{\rm SF\ now}$ allows us to maintain a SF efficiency close to $\epsilon_{\rm SF}$ during the entire simulation.
Once the total mass in new stars has been decided, we draw stars from the IMF until the required mass has been reached. A safe check is imposed such that if $M_{\rm new\ stars}>M_{\rm gas,2}$, then star formation is postponed to the next timestep.